# AI-Assisted Human Evaluation of Machine Translation

### v1

@@ Abstract
Annually, research teams spend large amounts of money to evaluate the quality of machine translation systems (WMT, Kocmi et al., 2023, inter alia). This is expensive because it requires detailed human labor. The recently proposed annotation protocol, Error Span Annotation (ESA), has annotators marking erroneous parts of the translation. In our work, we help the annotators by pre-filling the span annotations with automatic quality estimation. With AI assistance, we obtain more detailed annotations while cutting down the time per span annotation by half (71s/error span $\rightarrow$ 31s/error span). The biggest advantage of ESAAI protocol is an accurate priming of annotators (pre-filled error spans) before they assign the final score as opposed to starting from scratch. In addition, the annotation budget can be reduced by up to 24% with filtering of examples that the AI deems to be very likely to be correct.
@@ 1 Introduction
The quality of machine translation (MT) systems is periodically evaluated by academic and industry teams to measure progress and inform deployment decisions. This undertaking at scale, such as the WMT campaigns (Kocmi et al., 2022, 2023, inter alia), is extremely expensive, when requiring high annotation quality. Despite recent advancements in automated metric design (Freitag et al., 2023), these metrics remain misaligned with the ideal measure of text quality and human evaluation remains the most accurate and reliable standard.
Human evaluation protocols range from ranking different system outputs against each other (Novikova et al., 2018), to predicting scores (direct assessment, DA, Graham et al., 2015), or predicting specific error spans, types, and their severities (Multidimensional Quality Metrics, MQM, Lommel et al., 2014; Freitag et al., 2021). Kocmi et al. (2024) simplified this protocol into Error Span Annotation (ESA), which focuses only on the error span severities but not the actual error types, and is thus faster. One of the problems of the existing annotation protocols is their either very high cost, or low quality. In this work we aim to make the MT evaluation process with ESA less expensive.
@@
We pose that human evaluation of MT can benefit from AI assistance. Despite the risk of automatic bias, human-AI collaboration can be faster and more accurate than human or AI alone (Bondi et al., 2022). Thus, instead of showing annotators just the source and the system translation, we pre-fill the translation with error annotations from an AI system (Figure 1 bottom). This setup, which we call ESAAI, is enabled by the advancements in quality estimation systems (Guerreiro et al., 2023; Fernandes et al., 2023; Kocmi and Federmann, 2023), which provide accurate initial error spans. The main advantage of ESAAI comes from priming the user with possible translation errors.
ESAAI yields 1.6 error spans per translation segment in contrast to 0.5 for human-only ESA. While the overall ESAAI annotation time is slightly lower to that of ESA (58$\rightarrow$52s/segment), ESAAI halves the time per span annotation (71$\rightarrow$31s/error span). In majority of cases where the AI did not predict any errors, the annotators did not add any new error span. We find that we can pre-filter such examples from the evaluation, save up to 24% of the budget, and the evaluation results will be almost identical. In addition, because of the unified priming, the annotators also become more self-consistent and have higher inter-annotator agreement, suggesting higher annotation quality.
@@ 2 Related Work
One of the goals of MT evaluation is to compare systems to inform decisions (e.g. which system to deploy). Reference-based metrics compare the system translation to gold human translation, which can introduce evaluation bias (Freitag et al., 2020, 2023; Zouhar and Bojar, 2024). Reference-less approaches, known as quality estimation, do not have this problem, but they do not always correspond to human perception of quality (Freitag et al., 2023; Zouhar et al., 2024; Falcão et al., 2024) because the task is more complex. In higher-stakes settings, human annotators are employed to judge the translation quality.
The simplest option for human evaluation is to show the source and the translation and ask the annotators to give a number from 0 to 100, DA (Graham et al., 2015; Kocmi et al., 2022). and has the issue of low reliability and agreement. To make the annotations more objective, we can ask the annotators to mark specific errors in the translation (Multidimensional Quality Metrics, MQM, Lommel et al., 2014; Freitag et al., 2021). The marking is done based on their severity (e.g. minor or major) but also type (e.g. “inconsistent terminology”). This requires well-trained annotators and is thus expensive. In addition, this protocol does not yield scores, but only error spans, which are turned into the final score with a handcrafted formula. [^1] With some exceptions, the score computation from spans is a sum across all errors with -1 for minor and -5 for major.
To simplify this process and align it with the goal of objective translation quality estimation, Kocmi et al. (2024) proposed ESA, which uses non-experts and asks them to provide only the error severity (not its type) and a final translation score. This combines both approaches in that the annotators are primed with their marked errors to provide high quality final scoring. The modalities are depicted in App. Figure 6.
@@
Prior work shows that annotators can benefit from AI assistance (Devarajan et al., 2023; Pavoni et al., 2022). However, the use of AI in evaluation is not straightforward because the AI might bias the user or induce over-reliance (Buçinca et al., 2021). Human annotators usually have a financial incentive to optimize their work. Veselovsky et al. (2023) showed, that up to 46% of annotators did use LLMs for abstractive summarization. Including AI assistance in the annotation directly could decrease the use of undisclosed tools. We do so with quality estimation (QE/AI) systems that mark error spans in the output. The most popular QE systems are xCOMET (Guerreiro et al., 2023), AutoMQM (Fernandes et al., 2023) and GEMBA (Kocmi and Federmann, 2023). The QE system is not always correct, but the output is vetted by a human annotator. The QE thus still offloads some of the work that a human would do and better primes the annotators for evaluation.
3 Setup
@@
We implement our study in Appraise (Federmann, 2018) and use GEMBA, a GPT-based quality estimation system. We adapt the ESA protocol, where errors are marked and annotated as either minor or major.222 Minor: style/grammar/lexical choice could be better;
Major: changes meaning, lowers usability. See Appendix C.
The initial error markings are done by the AI and then post-edited by annotators. Afterwards, they manually assign a final score on the scale from 0 to 100% (not with AI). The error annotation part thus works as priming of the annotators in giving more accurate scores. The complete pipeline is shown in Figure 1 top. We also run the ESAAI setup twice with a different set of annotators to be able to determine the inter-annotator agreement and annotation stability.
@@
We first examine the high-level distribution of the collected data in Table 1. For ESAAI, the total number of reported error spans is three times higher than for ESA, which is due to the high number of suggested annotations by GEMBA. The split between minor and major errors is similar, though ESAAI annotators prefer major errors as opposed to ESA, even slightly more than those produced by GEMBA. Finally, the overall translation score is lower for ESAAI as opposed to just ESA. This is potentially caused by the priming effect of initially annotated error spans by GEMBA which highlight the negative aspects of the translation.
Protocol/method #errors %minor/%major Score
@@
GEMBA (AI only) 1.51 55% / 45% $\times$
Table 1: Average number of error spans and scores across ESA, ESAAI, and GEMBA.
4 Analysis
To evaluate the new ESAAI annotation pipeline, we consider two main aspects: (1) the annotation process, including its reliability and human effort, and (2) its usefulness for MT system comparisons.
@@ 4.1 ESAAI Evaluation Process
Not all post-editing operations are of equal value. For example, moving the error span by a few characters to the left is less important than adding a new error span for a missing translation. We point out two post-editing types: (1) changing the error span severity, and (2) editing the error span boundaries (App. Table 5). In 11% of cases, the users only changed the severity. This is important from the workflow perspective, because it only requires clicking the error span. In many cases, the error span was only moved. Time-wise this is more expensive because it requires the original error span to be removed and a new one created in its place. This operation can be skipped because it does not contribute to the ESA score. Therefore, the annotators could be instructed more specifically to not try to post-edit errors as long as they are approximately correct. See App. Table 9 for post-editing types.
Figure 2: Number of removed/kept/added error spans from GEMBA with respect to annotator progress.
Do annotators blindly accept AI hints? Gradual overreliance (Holford, 2022) is a type of automation bias which arises through repetition of non-problematic examples. Especially when there are no repercussions, the annotator might be tempted to only confirm the AI suggestion without actually doing any post-editing work. We first examine this through the perspective of changing in annotator’s behaviour thorough the annotation. In Figure 2 we show that the annotators make the same number of edits at the beginning as at the end, therefore excluding the possibility of automation bias.
@@
We use attention checks, where the translation is malformed but GEMBA does not show an error (App. Figure 8). The range for passing the attention checks for both protocols is the same—around 75% (App. Table 6). ESAAI is at a disadvantage because GEMBA spans were showing that the perturbed span was correct (i.e. no error spans). Therefore, the pertrubed examples were even more out-of-distribution and attention in-distribution is likely higher.
@@
Do AI mistakes affect annotators? Showing incongruent examples, where AI predictions are clearly wrong, has the potential of reducing the user’s trust in the AI and subsequent collaborating (Dhuliawala et al., 2023). Such examples are our attention checks. 84% and 73% of GEMBA-suggested spans are accepted for the document directly before and after the perturbed one. This is a slight decrease in trust, but it does not render the collaboration ineffective.
@@
How long do annotations take? One of the motivations of the AI-assisted setup is speeding up the annotations and leading to lower costs. The variance in individual annotator time can be explained by how much they post-edited the GEMBA error span annotation (see Figure 3). Per segment, ESAAI annotators required 52s while ESA required 58s, which is comparable. The time is 71s per single error span for ESA but 31s per single error span for ESAAI, making the latter more efficient in detailed annotation. In addition, the annotators get faster as the annotation progresses (Appendix B).
@@
Figure 3: Annotation actions (remove/keep/add an error span) and time per segment. Each dot is an annotator.
@@
Do annotators agree? For a robust and objective annotation protocol, the scores by two independent annotators should be similar. To test this, we ran the annotations again with different annotators. App. Table 7 shows that ESAAI has a much larger inter-annotator agreement. For the MQM-like score computation, this is due to the bias by the pre-filled error spans. Still, the agreement is much higher also for the direct scoring, likely due to the priming. This is consistent with much higher ESAAI intra-annotator agreement (App. Figure 10), i.e. how much annotators agree with themselves.

### v2

@@ Abstract
Annually, research teams spend large amounts of money to evaluate the quality of machine translation systems (WMT, Kocmi et al., 2023, inter alia). This is expensive because it requires a lot of expert human labor. The recently adopted annotation protocol, Error Span Annotation (ESA), has annotators marking erroneous parts of the translation and then assigning a final score. A lot of the annotator time is spent on scanning the translation for possible errors. In our work, we help the annotators by pre-filling the error annotations with recall-oriented automatic quality estimation. With this AI assistance, we obtain annotations at the same quality level while cutting down the time per span annotation by half (71s/error span $\rightarrow$ 31s/error span). The biggest advantage of ESAAI protocol is an accurate priming of annotators (pre-filled error spans) before they assign the final score. This also alleviates a potential automation bias, which we confirm to be low. In addition, the annotation budget can be reduced by almost 25% with filtering of examples that the AI deems to be very likely to be correct.
@@ 1 Introduction
The quality of machine translation (MT) systems is periodically evaluated by academic and industry teams to measure progress and inform product deployment decisions. This undertaking at scale, such as the WMT campaigns (Kocmi et al., 2022, 2023, inter alia), is extremely expensive. For high-quality systems, expensive high annotation quality is increasingly required to distinguish which system is truly better. Despite recent advancements in automated metric design (Freitag et al., 2023), the metrics remain misaligned with the ideal measure of text quality and human evaluation remains the most accurate, reliable, and ultimate standard.
@@
Human evaluation protocols range from ranking different system outputs against each other (Novikova et al., 2018), to assigning scores (direct assessment, DA, Graham et al., 2015), or marking specific error spans, types, and their severities (Multidimensional Quality Metrics, MQM, Lommel et al., 2014; Freitag et al., 2021). Kocmi et al. (2024) simplified the MQM protocol into Error Span Annotation (ESA), which focuses on the error span severities and not the actual error types. At the end, the annotators additionally assign a final score to the translation. The ESA protocol thus combines the objective diagnostics qualities of MQM (error spans), with the speed and evaluation focus of DA (scoring). One of the problems of all the existing annotation protocols is still their either very high cost, or low quality. In this work we aim to make the MT evaluation process with ESA less expensive.
In the past, human translation has benefited by being reframed as post-editing of machine translation (Zouhar et al., 2021). In this work, we propose that human evaluation of MT can benefit from AI assistance in a similar way. Despite the risk of automatic bias (blindly accepting AI suggestions), human-AI collaboration can be faster and more accurate than human or AI alone (Bondi et al., 2022). Thus, instead of showing annotators just the source and the system translation, we pre-fill the translation with error annotations from an AI system (Figure 1 bottom). This is motivated by a lot of human labor being spent on finding possible errors and we fill this with recall-focused quality estimation system that produces error annotations. The users still edit the error spans but now spend less effort on scanning the translation for possible errors. This setup, which we call ESAAI, is enabled by the advancements in quality estimation systems (Guerreiro et al., 2023; Fernandes et al., 2023; Kocmi and Federmann, 2023), which provide accurate initial error spans. The advantage of ESAAI comes not just from the error span suggestions but also from priming the user with possible translation errors before they assign the final score.
To test our setup, we conduct an translation evaluation annotation campaign contrastively with the ESA and ESAAI protocols. We compare these protocols on speed, inter-/intra-annotator agreement, quality control success, but also on a new meta-metric subset consistency accuracy.
Findings.
ESAAI yields 1.6 error spans per translation segment in contrast to 0.5 for human-only ESA. While the overall ESAAI annotation time is only slightly lower to that of ESA (58$\rightarrow$52s/segment), ESAAI halves the time per error span annotation (71$\rightarrow$31s/error span). This is because the output of ESAAI is more than three times of error spans per segment than from ESA.
In majority of cases where the AI did not predict any errors, the annotators did not add any new error span, confirming high-recall of the AI. We also find that we can pre-filter such examples from the evaluation, save up to 24% of the budget, and the evaluation results will be almost identical. In addition, because of the unified priming, the annotators also become more self-consistent and have higher inter-annotator agreement, suggesting higher annotation quality. Ultimately, this allows for lower number of annotators required to arrive at the same system ranking (high subset consistency accuracy).
Figure 2: Overview of inputs and outputs of various machine translation evaluation approaches. Quality estimation (QE) is automated and produces for each segment either a single score or a list of errors. DA+SQM, MQM, ESA and ESAAI are human annotation protocols. ESAAI (this paper) is semi-automated and happens in two-steps: quality estimation pre-annotation and human annotation.
@@ 2 Related Work
One of the goals of MT evaluation is to compare systems to inform decisions such as which system to deploy or which machine learning method works the best. There are two ways in which to evaluate translation quality: with automatic metrics or with human labor. Reference-based metrics compare the system translation to gold human translation. They do not always correspond to human perception of quality and can also introduce evaluation bias (Freitag et al., 2020, 2023; Zouhar and Bojar, 2024). Reference-less approaches, known as quality estimation (QE), do not have the reference bias problem, but also do not always correlate with human judgement (Freitag et al., 2023; Zouhar et al., 2024; Falcão et al., 2024) because the task is more difficult. In higher-stakes settings, human annotators are thus always employed to reliably judge the translation quality.
We depict the various human evaluation protocols in Figure 2. The simplest option for human evaluation is to show annotators the source and the translation and ask them to give a number from 0 to 100, DA (Graham et al., 2015; Kocmi et al., 2022). but has the issue of low reliability and agreement. To make the annotations more objective, one can ask the annotators to mark specific errors in the translation (Multidimensional Quality Metrics, MQM, Lommel et al., 2014; Freitag et al., 2021). The marking is done based on their severity (e.g. minor or major) but also type (e.g. “inconsistent terminology”). This requires well-trained professional annotators and is thus expensive. In addition, this protocol does not yield scores, but only error spans, which are turned into the final score with a handcrafted formula that can introduce additional problems. With some exceptions, the score computation from spans is a sum across all errors with -1 for minor and -5 for major.
Error Span Annotation (ESA).
To simplify this process and align it with the goal of objective translation quality estimation, Kocmi et al. (2024) proposed ESA, which asks annotators to provide only the error severity (not its type) but also a final translation score. Because the error type is not required, this makes the whole annotation faster and also non-experts can be employed because the annotators do not need to know the error type ontology. This combines both DA and MQM in that the annotators are primed with their marked errors to provide high quality final scoring. The modalities are depicted in the penultimate row in Figure 2.
@@
Prior work shows that annotators can benefit from AI assistance (Devarajan et al., 2023; Pavoni et al., 2022). However, the use of AI in evaluation is not straightforward because the AI might bias the user or induce over-reliance (Buçinca et al., 2021) where the annotator mindlessly accepts AI suggestions to reduce their labor. This can happen because human annotators usually have a financial incentive to optimize their work. In addition, Veselovsky et al. (2023) showed, that up to 46% of annotators did use LLMs for abstractive summarization. Including AI assistance in the annotation directly could therefore decrease the use of undisclosed tools. See Figure 2 for a depiction how this combines quality estimation pipeline and the ESA protocol.
Quality estimation.
We do include the AI assistance with quality estimation (QE) systems that mark error spans in the output. Specifically, given the source and only the translation (i.e. not the reference), it produces error span annotations (see Figure 2). Because it is not dependent on the reference, it can also be used in more setups e.g. where the reference is only being created through this process. Despite the history of quality estimation, such explainable QE has become popular only recently (Fomicheva et al., 2021). The most popular QE systems are xCOMET (Guerreiro et al., 2023), AutoMQM (Fernandes et al., 2023) and GEMBA (Kocmi and Federmann, 2023).
GEMBA, the QE system that we use, is based on prompting a GPT-4 model, and therefore easily adaptable to new scenarios. With a one-shot example in the target language, the model is prompted to provide a list of MQM-like errors. We only use the error spans and severities from this model. See the full prompts in Appendix C.
The QE system is not always correct, but the output is vetted by a human annotator. Compared to humans, the QE system is recall-focused, thus erring on the side of highlighting spans that are not erroneous. Removing false positives is easier and faster for a human annotator than to scan the whole translation for false negatives. The QE thus still offloads some of the work that a human would do and better primes the annotators for final score evaluation.
3 Machine Translation evaluation withHuman-AI collaboration
With improved machine translation systems, distinguishing which one is the best becomes increasingly difficult, requiring experts to annotate more and more samples. Some parts of the expert human evaluation do not require full attention or can be automated. With this, we reframe human evaluation as a computer-assisted annotation task to allow for future-proof scaling where competing systems’ quality requires more evaluated samples.
We now describe the technical details needed for exact study replication.
@@ 3 Machine Translation evaluation withHuman-AI collaboration
We implement our study in Appraise (Federmann, 2018) and use GEMBA, a GPT-based quality estimation system. We adapt the Error Span Annotation (ESA) protocol (Section 2), where errors are marked on character level and annotated as either minor or major.111 Minor: style/grammar/lexical choice could be better;
Major: changes meaning, lowers usability. See Appendix B.
The initial error markings are done by the AI and then post-edited by annotators. Afterwards, the annotators manually assign a final score on the scale from 0% to 100% (not with AI) ranging from “no meaning preserved” to “perfect”. See interface screenshot in App. Figure 11 and guidelines in Appendix B. The error annotation part thus works as priming of the annotators in giving more accurate scores. The complete pipeline is shown in Figure 1 top. We run the ESAAI setup twice with a different set of annotators to be able to determine the inter-annotator agreement and annotation stability. We finally request about 30% of annotators to re-do their work two months later to estimate the intra-annotator agreement, also known as self-consistency. We hire 21 annotators that are professional translators and native in the target language, German.
Dataset and collected data.
We use the data of WMT23 Metrics Shared Task (Freitag et al., 2023) which has been annotated with MQM and ESA. WMT23 is the latest iteration of the Conference on Machine Translation which asks research and industry teams to submit their machine translation systems. These systems are then evaluated with human experts to determine the final system ranking. This ranking is useful, amongst other things, for measuring research trends and overall improvements. However, because the submitted systems are state-of-the-art and recently close to human quality, arriving at the ranking requires more and more annotations. This motivates finding annotation protocols that speed up the annotations without sacrificing quality.
@@
4 Analyzing ESAAI Efficacy
To evaluate the new ESAAI annotation pipeline, we consider two main aspects: (1) the annotation process, including its reliability and human effort, and (2) its usefulness for machine translation systems comparison and costs.
#errors Minor/Major Score
@@ 4 Analyzing ESAAI Efficacy
QE (automated) 1.51 55% / 45% $\times$
Table 1: Average number of error spans and scores across ESA, ESAAI, and the QE system (automated). Because of the pre-annotations, the output of ESAAI is much more errors than ESA alone.
Operation Frequency
Severity change 12.0%
                Increase severity 60.0%
                Decrease severity 40.0%
Move span $\leq$5 13.1%
Move span $\leq$10 17.2%
Move span $\leq$20 23.3%
Resize      Increase error span size 21.5%
                Decrease error span size 78.5%
Table 2: Distribution of two ESAAI post-editing types: changing the severity, and moving the error span. A span is considered to be moved if the distance between old and new endpoints is at most 5, 10, or 20 characters. Many of the QE errors are only misplaced or have the wrong severity. See specific cases in Table 3.
@@ 4.1 ESAAI Evaluation Process
Collected data distribution.
We first examine the high-level distribution of the collected data in Table 1. For ESAAI, the total number of reported error spans is three times higher than for ESA, which is due to the high number of suggested annotations by the QE system. The split between minor and major errors is similar, though ESAAI annotators prefer major errors as opposed to ESA, even slightly more than those produced by the QE system. Finally, the overall translation score is lower for ESAAI as opposed to just ESA. This is potentially caused by the priming effect of initially annotated error spans by the QE which highlight the negative aspects of the translation.
QE Die Physik ist schrecklich und die Leute, die das Spiel entwickelt haben, werden nichts dagegen tun
Decrease severity Source Will not buy Mr. Coffee again
QE Ich kaufe Mr. Kaffee nicht mehr.
ESAAI Ich kaufe Mr. Kaffee nicht mehr.
QE Aber ich hasse Kunst und Literatur, und meine Schulgeschichte bestätigt es.
Resize Source [$\ldots$] I’m not sure if that would work for this.
QE [$\ldots$] ich bin mir nicht sicher, ob das für diesen Zweck funktionieren würde.
{example}
What post-edits do annotators make?
Not all post-editing operations are of equal value. For example, moving the error span by a few characters to the left is less important than adding a new error span for a missing translation. We point out two post-editing types: (1) changing the error span severity, and (2) editing the error span boundaries (Table 2). In 11% of cases, the users only changed the severity. This is important from the workflow perspective, because it only requires clicking the error span. In many cases, the error span was only moved. Time-wise this is more expensive because it requires the original error span to be removed and a new one created in its place. This operation can be skipped because it does not contribute to the ESA score. Therefore, the annotators could be instructed more specifically to not try to post-edit errors as long as they are approximately correct. See Table 3 for post-editing types.
Figure 3: Number of removed/kept/added error spans from the QE system with respect to annotator progress. The amount and type of work remains constant.
Do annotators blindly accept AI hints?
Gradual overreliance (Holford, 2022) is a type of habituation or automation bias which arises through repetition of non-problematic examples, such as cancer diagnosis which is dominantly negative. Especially when there are no repercussions, the annotator might be tempted to only confirm the AI suggestion without actually doing any post-editing work. As a result, they would either confirm a span that is not an error or miss part of the translation that is not highlighted by AI but is, in fact, erroneous. We first examine this through the perspective of changes in annotator’s behaviour thorough the annotation. In Figure 3 we show that the annotators make the same number of edits at the beginning as at the end, therefore excluding the possibility of learned automation bias. Next, we examine if annotators are not overreliant on the AI already from the beginning.
@@
Attention checks, stimuli where we apriori know the correct annotation, are a mechanism to verify that the annotator does the expected job. We use attention checks where the translation is intentionally malformed but the QE system does not show an error (Figure 4). Per 100 segments, there are 12 attention checks in total, each with one perturbed span, and each annotator sees both the attention check and the translation original, randomly shuffled. This way, we can compare the annotator’s score between the perturbed and non-perturbed versions. The range for passing the attention checks (on score, error count, or error highlight level) for ESA is 65% and for ESAAI 69% (Table 4). This is despite ESAAI being at a disadvantage because the segments, as in Figure 4, contain errors that are strictly not highlighted by the QE system. Therefore, the pertrubed examples were even more out-of-distribution and the attention of the annotators in-distribution is likely higher.
SRC:   Sie haben gestern das Treffen wieder verschoben.
TGT:   He postponed the meeting again yesterday.
[…]
{example}
[…]
Original Perturbed OK
ESA Score 79.5 52.6 86%
Span count 0.85 1.86 54%
Perturbation marked 56%
ESAAI Score 75.8 52.6 76%
Span count 2.19 4.48 61%
Perturbation marked 71%
Do AI mistakes affect annotators?
@@
Showing incongruent examples, where AI predictions are clearly wrong, has the potential of reducing the user’s trust in the AI and their subsequent collaborating (Dhuliawala et al., 2023). In our case, such examples are the attention checks where the AI intentionally misses the perturbed part. To measure the effect of this, we look at documents directly preceding and following the attention check. 84% of AI-suggested spans are accepted for the document directly before. In contrast, documents directly after the attention check has only 73% of acceptance of AI-suggested spans. This is a slight decrease in trust, but it does not render the collaboration ineffective. It also shows that the annotators are sensitive to possible AI mistakes.
How long do annotations take?
@@
One of the motivations of the AI-assisted setup is speeding up the annotations and lowering the costs. The variance in individual annotator time can be explained by how much they post-edited the QE system’s error span annotation (see Figure 5). In this aspect, ESAAI has more variance than ESA, but also can be more controlled and constrained by instructing the annotators what the expected post-editing level is. Per segment, ESAAI annotators required 52s while ESA required 58s. In addition, the time is 71s per single error span for ESA but 31s per single error span for ESAAI, making the latter more efficient in detailed annotation.
@@
Figure 5: Annotation actions (remove/keep/add an error span) and time per segment. Each dot and bar is an annotator (sorted by time).
Do annotators agree?
@@
For a robust and objective annotation protocol, the scores by two independent annotators should be similar and not subjective. To test this, we ran the annotations again with different annotators. Table 5 shows that ESAAI has a much larger inter-annotator agreement. For the MQM-like score computation from spans, this is due to the bias by the pre-filled error spans. Still, the agreement is much higher also for the direct scoring, likely due to the unified priming of the annotators. This is consistent with much higher ESAAI intra-annotator agreement, i.e. how much annotators agree with themselves.
inter-annotator intra-annotator
Scoring ESA ESAAI ESA ESAAI
direct score 0.376 0.533 0.222 0.486
from spans 0.327 0.671 0.282 0.689
Table 5: Inter-annotator and intra-annotator agreement with direct scores and scores computed from error spans with MQM formula, as measured with Spearman correlation. ESAAI from spans have the highest inter-annotator agreement, which is however caused by the the QE system’s pre-filling. Still, the scores from ESAAI, solely by humans, have the highest inter-annotator and intra-annotator agreement. See visualization in App. LABEL:{fig:intrainter_annotator_agreement}.
Do annotators become faster?
With most annotations task, the annotators learn to be faster. Even though the speedup happens thorough the whole annotation, it is mostly present in the first 15% of segments (green box in Figure 6). The ESAAI annotators get 1.87s faster with every segment, which is comparable to ESA. This effect is present despite the ESA annotators being at an advantage because there was more ESAAI annotators and they thus each individually processed fewer segments. In addition, users in the post-editing task are more consistent. For ESA, the user’s absolute deviation from their personal average is 43.3s, while for ESAAI this is only 32.1s. This makes the human effort more consistent and predictable but also shows that the nature of the annotation task changes.
Why do some segments take longer than others?
Being able to predict the expected annotation time for a segment can lead to a more efficient distribution and planning, for example when selecting which segments to annotate at all. We examine the correlations between features and the segment-level time in Table 6. The number of words in the translation, together with the number of error spans is a strong predictor of annotation time. For MQM this is the highest, which can be explained by each error span requiring the most work in the MQM annotation scheme because the annotators have to also assign the error type. The longer the whole document (number of surrounding translation segments), the lower the annotation time, which is likely due to shared context so the annotator does not have to switch between domains and contexts. Contrastively, the ESAAI annotators are slightly less affected by the translation length in contrast to ESA because the error spans are pre-highlighted.
MQM ESA ESAAI
Progress -0.12 -0.13 -0.13
Translation word count 0.30 0.19 0.16
QE error spans 0.12 0.07 0.12
Error spans 0.06 0.04 0.12
Score -0.07 -0.03 -0.08
Document size -0.14 -0.17 -0.17